OpenStudy (khariangel290):

For a circle of radius 3 feet, find the arc length s subtended by a central angle of 57 degrees. the wanted it in radians but I keep getting the wrong answer.

OpenStudy (prathamesh_m):

To convert degrees to radians you multiply it by pi/180

OpenStudy (prathamesh_m):

Then multiply that with the radius to get the arc length.

OpenStudy (sshayer):

\[\theta=\frac{ l }{ r },l=r \theta \] where l is the length of arc. r=radius of circle \[\theta=\angle~\in~radians.\]